OpenStudy (anonymous):

The average of A and B is 6. The average of X,Y, and Z is 16. What is the average of A,B,X,Y, and Z?

OpenStudy (danjs):

(A+B+X+y+z) / 5 = ?

OpenStudy (anonymous):

\[\frac{ A+B }{ 2 } = 6\] \[A+B = 6*2\] \[\frac{ X+Y+Z }{ 3 } = 16\] \[X+Y+Z = 16*3\] \[\frac{ A+B+X+Y+Z }{ 5 } = ?\] \[\frac{ (A+B) + (X+Y+Z) }{ 5 } = ?\] Put the value of A+B and X+Y+Z